Let A = {1,2,3,4,5,6} and let p1 = (3,6,2) and p2 = (5, 1, 4) be permutations of A.

(a) Compute p1 ◦ p2 and write the result as a product of cycles and as the product of transpositions.

(b) Compute p−1 ◦ p−1. 


1
Expert's answer
2022-01-31T15:51:08-0500

"p_1=(362), p_2=(514)\\
p_1 \circ p_2 =(362) \circ (514)=\begin{pmatrix}
 1 & 2&3&4&5&6 \\
 4 & 3 &6&5&1&2
\end{pmatrix}=(145)(236)"


b.)

"p_1^{-1}=(362)^{-1}=(263)\\
p_2^{-1}=(514)^{-1}=(154)\\
p_1^{-1} \circ p_2^{-1}=(263)\circ(154)=\begin{pmatrix}
 1&2&3&4&5&6 \\
 5&6&2&1&4&3
\end{pmatrix}=(154)(263)"